# Structure-preserving long-time simulations of turbulence in magnetized ideal fluids

### v1

@@ 1. Introduction
The Lie–Poisson structure and its associated Hamiltonian dynamics impose geometrical properties in phase space that affect the long-time statistical behaviour (an example is the inverse energy cascade in 2-D hydrodynamics). Thus, in the discretization of Lie–Poisson systems it is natural to aim for numerical schemes that preserve as much as possible of the structure, to retain the right statistical behaviour. In particular, the Lie–Poisson structure imposes invariant submanifolds in phase space called co-adjoint orbits, and the dynamics is confined to these orbits. However, traditional numerical schemes, such as finite element, finite volume, or spectral methods, are typically designed to provide high accuracy in terms of minimizing the local error, but they fail to preserve the geometric structure, particularly the confinement to co-adjoint orbits. For long-time simulations of chaotic systems, such as those discussed in this paper, high accuracy is less important, because small perturbations grow exponentially (or faster). Instead, to retain the statistically correct qualitative behaviour, it is the phase space structure (such as co-adjoint orbits) that is important to capture numerically. Indeed, it is the geometrical structure, not the local dynamics, that is responsible for the formation of large scale structures in the long-time behaviour, such as the presence of the inverse cascade of energy in the case of hydrodynamical turbulence, and of the mean square of magnetic potential in MHD turbulence. One way to see this is via statistical mechanics, which is surprisingly useful for predicting the long-time behaviour, and which disregards the local dynamics altogether, but is directly affected by the geometric structure (in particular, the conservation laws and the symplectic nature of the phase flow). Geometric numerical methods, i.e., numerical discretizations that preserve the underlying phase space geometry, offer a middle ground: they take the geometry into account, while still imitating the local dynamics (albeit with lower accuracy than traditional high-order methods).
@@
Another geometric discretization algorithm was developed by the authors (Modin and Roop, 2025). The key idea is to replace the infinite-dimensional Lie–Poisson system with a finite-dimensional matrix analogue, following the approach of Zeitlin (2004, 1991, 2005) who, based on geometric quantization (Hoppe, 1982, 1989; Bordemann et al., 1991, 1994; Hoppe and Yau, 1998), developed matrix approximations for ideal fluids and RMHD on the flat torus (the doubly periodic square). The matrix dynamical system studied by Modin and Roop (2025) is itself a Lie–Poisson flow for a modified Lie–Poisson structure and retains (discrete versions of) the conservation laws, such as Casimir functions (magnetic helicity and cross-helicity) and energy. Furthermore, the modified Lie–Poisson structure, induced by the matrix equations, converge, along with the Casimirs, to the infinite-dimensional Lie–Poisson structure as the matrix dimension increases. As a full spatio-temporal discretization, the method gives a “magnetic” extension of the structure preserving matrix hydrodynamics approach to ideal fluids on the sphere (Modin and Viviani, 2020a, b). Although the method is specifically focused on the case of the sphere, the matrix equations for spherical RMHD are modelled on the same Lie algebra as those on the flat torus, so the time integrator can also be applied to RMHD on the torus. The sphere is, however, the preferred choice for simulation of Zeitlin’s equations because it fully preserves the invariance under the group of isometries, contrary to the corresponding matrix model for the flat torus (see the paper by Modin and Viviani (2024) for details).
The aim of the present paper is (i) to numerically investigate basic properties of long-time turbulence of magnetic fluids on the sphere as modelled by the RMHD equations, Hazeltine’s model, and the CHM model, and (ii) to compare the resulting long-time behaviour between the models. A motivation for our study comes from previous simulation results on hydrodynamic turbulence on the sphere (Cifani et al., 2023; Modin and Viviani, 2020b), which demonstrate the formation of large-scale condensates. Do formations like those appear also in the various models for reduced magnetized fluids?
@@
Acknowledgements. This work was supported by the Swedish Research Council (grant number 2022-03453), the Knut and Alice Wallenberg Foundation (grant number WAF2019.0201), and the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. The work of M.R. is supported by grants from the Royal Swedish Academy of Sciences (MA2024-0034, MG2024-0050). The computations were enabled by resources provided by Chalmers e-Commons at Chalmers. The data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. The authors would like to thank Darryl Holm and Philip Morrison for pointing us to Hazeltine’s model for magnetic fluids.
@@ 2.1. Reduced magnetohydrodynamics (RMHD)
These invariants are Casimirs, i.e., for any functional $\mathcal{J}(\omega,\theta)$, one has $\llbracket\mathcal{C}_{f},\mathcal{J}\rrbracket=\llbracket\mathcal{I}_{g},\mathcal{J}\rrbracket=0$. We observe that $g(\theta)=\theta$ corresponds to the cross-helicity invariant, and $f(\theta)=\theta^{2}$ corresponds to the mean-square magnetic potential, both critical in MHD turbulence. More generally, we refer to $\mathcal{C}_{f}$ as magnetic helicity and to $\mathcal{I}_{g}$ as cross-helicity. Of course, the Hamiltonian (2.2) is also conserved by the flow (2.1), but it is not a Casimir invariant.
@@ 2.2. Hazeltine’s model and the Charney-Hasegawa-Mima (CHM) model
$\begin{split}\llbracket F,G\rrbracket&=\int\limits_{M}\left[\omega\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\omega}\right\}+\theta\left(\left\{\frac{\delta F}{\delta\theta},\frac{\delta G}{\delta\omega}\right\}+\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\theta}\right\}\right)+\right.{}\\ &\left.\chi\left(\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\omega}\right\}+\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\chi}\right\}\right)+\chi\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\chi}\right\}+\theta\left(\left\{\frac{\delta F}{\delta\theta},\frac{\delta G}{\delta\chi}\right\}+\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\theta}\right\}\right)\right]\mathrm{d}x\mathrm{d}y,\end{split}$ (2.9)
@@ 3. Matrix discretizations
In this section, we give spatio-temporal discretizations for the systems (2.1) and (2.5) based on finite-dimensional, matrix versions of the Lie–Poisson structures discussed previously. The main quality of such discretizations is that they capture all the geometric structure in the problem, rather than focusing on local accuracy as traditional numerical discretizations do. The idea goes back to Zeitlin (Zeitlin, 1991, 2004, 2005) and is based on geometric quantization theory (Hoppe, 1982, 1989; Bordemann et al., 1991, 1994; Hoppe and Yau, 1998).
From here on, we fix the manifold $M$ to be the two-sphere, $M=S^{2}$. Although Zeitlin’s model was originally developed for the flat torus, it performs better on the sphere because of its consistency with the $SO(3)$ symmetry of the sphere (cf. (Modin and Viviani, 2024)). Albeit most MHD simulations in the literature are on the flat torus (doubly periodic square), we view the sphere as an equally suitable domain to investigate basic properties of MHD turbulence.
@@ 3.1. The RMHD–Zeitlin equations
To obtain the connection between matrices, such as $W$, and fields, such as $\omega$, we consider the eigen-matrices of the Hoppe–Yau Laplacian $\Delta_{N}$, which form a basis $T_{lm}^{N}$, $l=0,1,\ldots,N-1,\,m=-l,\ldots,l$ for $\mathfrak{su}(N)$ called matrix harmonics. Then, for $W\in\mathfrak{su}(N)$, one reconstructs the vorticity function $\omega\in C^{\infty}_{0}(S^{2})$, up to the first $N^{2}$ terms of its Fourier decomposition in the spherical harmonics basis $Y_{lm}$, by the correspondences $T_{lm}^{N}\sim Y_{lm}$. For more details on mapping between the vorticity matrix $W\in\mathfrak{su}(N)$ and the vorticity function $\omega\in C^{\infty}_{0}(S^{2})$, we refer to the publications (Modin and Viviani, 2024; Modin and Roop, 2025).
@@
Likewise, the discretized Casimirs invariants are
@@
As $N\to\infty$, the Hamiltonian (3.2) and the Casimirs (3.3) converge to their continuous counterparts (2.2) and (2.4) (for $f$ and $g$ taken as monomials, see (Modin and Roop, 2025) for the exact result).
@@
where $\tilde{P}=\Delta_{N}^{-1}\tilde{W}$, $\tilde{J}=\Delta_{N}\tilde{\Theta}$, and $\varepsilon=\delta t/\hbar$ for the physical time step length $\delta t>0$. The integrator (3.4) exactly [^1] Up to the machine precision of floating-point numbers when implemented on a computer. preserves the Casimirs (3.3) for any choice of functions $f$ and $g$, and nearly preserves the Hamiltonian (3.2), see (Modin and Roop, 2025).
@@ 3.2. The Hazeltine–Zeitlin and CHM–Zeitlin equations
In this case, Lie–Poisson time integration is obtained via the isospectral midpoint method (Modin and Viviani, 2020a)
@@ 4.1. Reduced magnetohydrodynamics (RMHD)
Figure 5. RMHD: Kinetic $E_{kin}$ and magnetic $E_{magn}$ energy distribution over time. The total energy $E_{magn}+E_{kin}$ is conserved. Magnetic and kinetic components of energy are redistributed with a tendency towards equipartition.
@@ 4.3. Hazeltine’s model
For the Hazeltine model simulation, we choose the length scale parameter $\alpha=1/2$, and prepare the initial vorticity distribution $\omega_{0}$ with vanishing total angular momentum. The motivation for such a choice comes from the long-time simulations of the incompressible Euler equations, where the dynamics settles at the four-blob configuration in the case of the vanishing total angular momentum; see, e.g., (Modin and Viviani, 2020b), where it is conjectured that vorticity mixing continues until the system reaches nearly integrable behaviour and remains close to it in the KAM sense. As in the previous simulations of the RMHD system, we choose initial condition with equal contribution to the total energy from all the fields. The spatial resolution is again $N=512$ and the total simulation time is set to $t=562$.
@@
We conclude this section with the dynamics of the density variation field $\chi(t)$. Although the $\chi(t)$ field fulfills the same equation as the field $\omega(t)$ in the RMHD model, one does not observe a vast amplification of $\chi(t)$ in the Hazeltine model, as seen from Figure 14. In the long-time dynamics, one observes circulations centred approximately at locations of the magnetic potential eddies.
@@
Figure 14. Hazeltine: Evolution of the density variation field $\chi(t)$.

### v2

@@ 1. Introduction
The Lie–Poisson structure and its associated Hamiltonian dynamics impose geometrical properties in phase space that affect the long-time statistical behaviour (an example is the inverse energy cascade in 2-D hydrodynamics). Thus, in the discretization of Lie–Poisson systems it is natural to aim for numerical schemes that preserve as much as possible of the structure, to retain the right statistical behaviour. In particular, the Lie–Poisson structure imposes invariant submanifolds in phase space called co-adjoint orbits, and the dynamics is confined to these orbits (for details on Lie–Poisson dynamics we refer to the monograph by Marsden and Ratiu (1999)). However, traditional numerical schemes, such as finite element, finite volume, or spectral methods, are typically designed to provide high accuracy in terms of minimizing the local error, but they fail to preserve the geometric structure, particularly the confinement to co-adjoint orbits. For long-time simulations of chaotic systems, such as those discussed in this paper, high accuracy is less important, because small perturbations grow exponentially (or faster). Instead, to retain the statistically correct qualitative behaviour, it is the phase space structure (such as co-adjoint orbits) that is important to capture numerically. Indeed, it is the geometrical structure, not the local dynamics, that is responsible for the formation of large scale structures in the long-time behaviour, such as the presence of the inverse cascade of energy in the case of hydrodynamical turbulence, and of the mean square of magnetic potential in MHD turbulence. One way to see this is via statistical mechanics, which is surprisingly useful for predicting the long-time behaviour, and which disregards the local dynamics altogether, but is directly affected by the geometric structure (in particular, the conservation laws and the symplectic nature of the phase flow). Geometric numerical methods, i.e., numerical discretizations that preserve the underlying phase space geometry offer a middle ground: they take the geometry into account, while still approximating the local dynamics. [^1] Although it is sometimes possible to construct geometric methods with as high order as traditional methods, it is usually more efficient to use low-order geometric methods, since the important property is conservation of phase space geometry, not high accuracy in tracking individual trajectories.
@@
Another geometric discretization algorithm was developed by the authors (Modin and Roop, 2025). The key idea is to replace the infinite-dimensional Lie–Poisson system with a finite-dimensional matrix analogue, following the approach of Zeitlin (2004, 1991, 2005) who, based on geometric quantization (Hoppe, 1982, 1989; Bordemann et al., 1991, 1994; Hoppe and Yau, 1998), developed matrix approximations for ideal fluids and RMHD on the flat torus (the doubly periodic square). The matrix dynamical system studied by Modin and Roop (2025) is itself a Lie–Poisson flow for a modified Lie–Poisson structure and retains (discrete versions of) the conservation laws, such as Casimir functions (magnetic helicity and cross-helicity) and energy. Furthermore, the modified Lie–Poisson structure, induced by the matrix equations, converge, along with the Casimirs, to the infinite-dimensional Lie–Poisson structure as the matrix dimension increases. As a full spatio-temporal discretization, the method gives a “magnetic” extension of the structure preserving matrix hydrodynamics approach to ideal fluids on the sphere (Modin and Viviani, 2020b, a). Although the method is specifically focused on the case of the sphere, the matrix equations for spherical RMHD are modelled on the same Lie algebra as those on the flat torus, so the time integrator can also be applied to RMHD on the torus. The sphere is, however, the preferred choice for simulation of Zeitlin’s equations because it fully preserves the invariance under the group of isometries, contrary to the corresponding matrix model for the flat torus (see the paper by Modin and Viviani (2024) for details).
The aim of the present paper is (i) to numerically investigate basic properties of long-time turbulence of magnetic fluids on the sphere as modelled by the RMHD equations, Hazeltine’s model, and the CHM model, and (ii) to compare the resulting long-time behaviour between the models. A motivation for our study comes from previous simulation results on hydrodynamic turbulence on the sphere (Cifani et al., 2023; Modin and Viviani, 2020a), which demonstrate the formation of large-scale condensates. Do formations like those appear also in the various models for reduced magnetized fluids?
@@
Acknowledgements. The authors would like to thank Darryl D. Holm and Philip J. Morrison for pointing us to Hazeltine’s model for magnetic fluids.
Funding. This work was supported by the Swedish Research Council (grant number 2022-03453); the Knut and Alice Wallenberg Foundation (grant number WAF2019.0201); the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine; and the Royal Swedish Academy of Sciences (M.R., grant numbers MA2024-0034, MG2024-0050, MA2025-0073). The computations were enabled by resources provided by Chalmers e-Commons at Chalmers. The data handling was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council (grant number 2022-06725).
@@ 2.1. Reduced magnetohydrodynamics (RMHD)
These invariants are Casimirs, i.e., for any functional $\mathcal{J}(\omega,\theta)$, one has $\llbracket\mathcal{C}_{f},\mathcal{J}\rrbracket=\llbracket\mathcal{I}_{g},\mathcal{J}\rrbracket=0$. We observe that $g(\theta)=\theta$ corresponds to the cross-helicity invariant, and $f(\theta)=\theta^{2}$ corresponds to the mean-square magnetic potential, both critical in MHD turbulence. More generally, we refer to $\mathcal{C}_{f}$ as magnetic helicity and to $\mathcal{I}_{g}$ as cross-helicity. Of course, the Hamiltonian (2.2) is also conserved by the flow (2.1), but it is not a Casimir invariant. Instead, its conservation follows from the anti-symmetry of the Lie–Poisson bracket.
@@ 2.2. Hazeltine’s model and the Charney-Hasegawa-Mima (CHM) model
$\begin{split}\llbracket F,G\rrbracket&=\int\limits_{M}\left[\omega\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\omega}\right\}+\theta\left(\left\{\frac{\delta F}{\delta\theta},\frac{\delta G}{\delta\omega}\right\}+\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\theta}\right\}\right)+\right.{}\\ &\left.\chi\left(\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\omega}\right\}+\left\{\frac{\delta F}{\delta\omega},\frac{\delta G}{\delta\chi}\right\}\right)+\chi\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\chi}\right\}+\right.{}\\ &\left.\theta\left(\left\{\frac{\delta F}{\delta\theta},\frac{\delta G}{\delta\chi}\right\}+\left\{\frac{\delta F}{\delta\chi},\frac{\delta G}{\delta\theta}\right\}\right)\right]\mathrm{d}x\mathrm{d}y,\end{split}$ (2.9)
@@ 3. Matrix discretizations
In this section, we give spatio-temporal discretizations for the systems (2.1) and (2.5) based on finite-dimensional, matrix versions of the Lie–Poisson structures discussed previously. The main quality of such discretizations is that they capture all the geometric structure in the problem, rather than focusing on local accuracy as traditional numerical discretizations do. The idea goes back to Zeitlin (1991, 2004, 2005) and is based on geometric quantization theory (Hoppe, 1982, 1989; Bordemann et al., 1991, 1994; Hoppe and Yau, 1998).
From here on, we fix the manifold $M$ to be the two-sphere, $M=S^{2}$. Although Zeitlin’s model was originally developed for the flat torus, it performs better on the sphere because of its consistency with the $SO(3)$ symmetry of the sphere (cf. Modin and Viviani (2024)). Albeit most MHD simulations in the literature are on the flat torus (doubly periodic square), we view the sphere as an equally suitable domain to investigate basic properties of MHD turbulence.
@@ 3.1. The RMHD–Zeitlin equations
To obtain the connection between matrices, such as $W$, and fields, such as $\omega$, we consider the eigen-matrices of the Hoppe–Yau Laplacian $\Delta_{N}$, which form a basis $T_{lm}^{N}$, $l=0,1,\ldots,N-1,\,m=-l,\ldots,l$ for $\mathfrak{su}(N)$ called matrix harmonics. Then, for $W\in\mathfrak{su}(N)$, one reconstructs the vorticity function $\omega\in C^{\infty}_{0}(S^{2})$, up to the first $N^{2}$ terms of its Fourier decomposition in the spherical harmonics basis $Y_{lm}$, by the correspondences $T_{lm}^{N}\sim Y_{lm}$. For more details on mapping between the vorticity matrix $W\in\mathfrak{su}(N)$ and the vorticity function $\omega\in C^{\infty}_{0}(S^{2})$, we refer to the publications by Modin and Viviani (2024); Modin and Roop (2025).
@@
Likewise, the discretized Casimir invariants are
@@
As $N\to\infty$, the Hamiltonian (3.2) and the Casimirs (3.3) converge to their continuous counterparts (2.2) and (2.4) (for $f$ and $g$ taken as monomials, see the paper by Modin and Roop (2025) for the exact result).
@@
where $\tilde{P}=\Delta_{N}^{-1}\tilde{W}$, $\tilde{J}=\Delta_{N}\tilde{\Theta}$, and $\varepsilon=\delta t/\hbar$ for the physical time step length $\delta t>0$. The integrator (3.4) exactly [^2] Up to the machine precision of floating-point numbers when implemented on a computer. preserves the Casimirs (3.3) for any choice of functions $f$ and $g$, and nearly preserves the Hamiltonian (3.2), see Modin and Roop (2025).
@@
Contrary to generic Lie–Poisson integrators, the method (3.4) is free of matrix exponentials and therefore scales efficiently as the matrix size grows. The complexity per time step is $O(N^{3})$, see Cifani et al. (2023).
@@ 3.2. The Hazeltine–Zeitlin and CHM–Zeitlin equations
In this case, Lie–Poisson time integration is obtained via the isospectral midpoint method (Modin and Viviani, 2020b)
@@ 4.1. Reduced magnetohydrodynamics (RMHD)
Figure 5. RMHD: (a) Total energy variation, and (b) kinetic $E_{kin}$ and magnetic $E_{magn}$ energy evolution over time. The total energy $E_{magn}+E_{kin}$ is conserved up to a relative error of about $10^{-5}$. The magnetic and kinetic energy components are redistributed with a tendency towards equipartition.
@@ 4.3. Hazeltine’s model
For the Hazeltine model simulation, we choose the length scale parameter $\alpha=1/2$, and prepare the initial vorticity distribution $\omega_{0}$ with vanishing total angular momentum. The motivation for such a choice comes from the long-time simulations of the incompressible Euler equations, where the dynamics settles at the four-blob configuration in the case of the vanishing total angular momentum; see, e.g., (Modin and Viviani, 2020a), where it is conjectured that vorticity mixing continues until the system reaches nearly integrable behaviour and remains close to it in the KAM sense. As in the previous simulations of the RMHD system, we choose initial condition with equal contribution to the total energy from all the fields. The spatial resolution is again $N=512$ and the total simulation time is set to $t=562$.
@@
We conclude this section with the dynamics of the density variation field $\chi(t)$. Although the $\chi(t)$ field fulfils the same equation as the field $\omega(t)$ in the RMHD model, one does not observe a vast amplification of $\chi(t)$ in the Hazeltine model, as seen from Figure 14. In the long-time dynamics, one observes circulations centred approximately at locations of the magnetic potential eddies.
@@
Figure 14. Hazeltine: Evolution of the density variation field $\chi(t)$. There are no clear large scale structures remaining at large times, although we observe some circulating filament-like structures centred about the locations of magnetic blobs in Figure 12.